(17.2b.3)

A 8.57×10-2-M aqueous solution of CH3NH2 (methylamine) has a pH of 11.8. Calculate the pH of a buffer solution that is 8.57×10-2 M in CH3NH2 and 0.156 M in CH3NH3+.
Answer #1

pH = 11.8

pOH = 2.2

[OH-] = 6.309 x 10^-3 M

CH3NH2    +    H2O     --------------> CH3NH3+ +     OH-

0.0857                                                     0                  0

0.0857-x                                                   x                  x

Kb = x^2 / 0.0857 - x

        = (6.309 x 10^-3)^2 / 0.0857 - 6.309 x 10^-3

Kb = 5.014 x 10^-4

pKb = 3.30

pKa = 10.70

pH = pKa + log [base / acid]

    = 10.70 + log [0.0857 / 0.156]

pH = 10.44